# An algorithm for forensic toolmark comparisons

### v2

@@ Abstract
Forensic toolmark analysis traditionally relies on subjective human judgment, leading to inconsistencies and inaccuracies. The multitude of variables, including angles and directions of mark generation, further complicates comparisons. To address this, we introduce a novel approach leveraging 3D data capturing toolmarks from various angles and directions. Through algorithmic training, we objectively compare toolmark signals, revealing clustering by tool rather than angle or direction. Our method utilizes similarity matrices and density plots to establish thresholds for classification, enabling the derivation of likelihood ratios for new mark pairs. With a cross-validated sensitivity of 98% and specificity of 96%, our approach enhances the reliability of toolmark analysis. While its applicability to diverse tools and factors warrants further exploration, this empirically trained, open-source solution offers forensic examiners a standardized means to objectively compare toolmarks, potentially curbing miscarriages of justice in the legal system.
@@ 1 Introduction
In this article, we present an open-source algorithm, an objective method for comparison of 3D scans of striated toolmarks. We made three contributions to the field. First, we generated three databases of toolmarks from consecutively manufactured flat-head unused screwdrivers: one to study the variability within and between tools at a fixed angle/direction, one to study the variability within and between angles of attack (80, 70, 60 with respect to the surface), and one to study the variability within and between directions of tool travel (pushing and pulling). We generate eight repetitions of the marks at each condition to study the variability within condition. Second, we used the GelSight portable handheld 3D scanner, which measures the 3D topography of a solid surface using elastomeric tactile sensor technology. This scanner is a portable and fast solution for collecting toolmarks at crime scenes without having to cast or remove the mark. Third, we ran a Louvain clustering algorithm on the three databases to determine whether the similarity within source (tool-side) was higher than the similarity between sources, at a fixed angle/direction, when varying angle, and when varying direction. Then we used what we learned from the clustering step to generate the same-source and different-source densities. We fit a Beta parametric distribution to these, and used an ensemble method (Guarch and Ommen, 2023), to generate likelihood ratios. Thus, a new pair or marks can be compared and a likelihood ratio can be produced using our method. Finally, we provide an R software package called tmaRks, an open-source implementation of this process, to make it openlyand freely accessible to forensic examiners and researchers.
@@ 2 Previous work
Baiker et al. (2014) has a number of interesting contributions. They use global cross correlation to compare marks made at five angles between 15-75 degrees from the normal. Their method has high discriminatory power even at different angles of attack. They find that their automated algorithm beats human examiners in terms of false positives and the humans beat the algorithm in terms of false negatives. They also show that relying on 3D data is better than relying on 2D data, and they find evidence that although their method was trained on certain screwdrivers, it could be used to compare other tools as well. Baiker et al. (2015) find that lead preserves fine details in the striations, that is advantageous to push the tool instead of pulling during toolmark creation for angles of attack above 45 degrees, and that toolmarks should be created as shallow as possible in the substrate material. Several other projects (Chumbley et al. (2009); Chumbley and Morris (2013); Grieve et al. (2014); Hadler and Morris (2018); Gambino et al. (2011); Petraco et al. (2012), to name a few) propose new machine learning and other methods to study 3D marks.
@@
Finally, there are several new publications in which researchers generate virtual toolmarks, which prevents the tool from being damaged while making test marks. Baiker et al. (2016) generate virtual toolmarks and compare these to empirical mark. They find that the virtual and empirical marks are very similar, and that they can recover the angle of attack from an experimental mark. They use wavelet analysis to make comparisons. Bachrach et al. (2013) also uses wavelet analysis to generate virtual bullet Land Engraved Area (LEA) signatures, and Ekstrand et al. (2014) developed a virtual toolmark generator that allows the user to specify the angle of attack and rotation. M.S. et al. (2015) predict angles from virtual marks.
@@ 3.1 Experimental design
We use a factorial design for the toolmarks generation, to allow for the study of the variability within source and between sources of marks, and how this changes as the conditions of of angle of attack and direction of mark generation change. Table 1 shows the combination of conditions for each. The factors are the angle and direction. We generate three sets of marks, which we call experiments, for a total of 560 marks. As can be seen in Table 1, we generate eight replicates per tool-side under the same conditions. Note that we used large screwdrivers for experiment 2, since varying angle required long screwdrivers to reach the lead plate.
Experiments 1, 2, and 3 allow us to study the variability between marks made by the same tool, since there are eight replicates made under each condition. The angles are 60, 70, and 80 degrees with respect to the lead surface. The directions are pushing and pulling, where pulling refers to a negative rake angle. For example, Figure 3 shows the marks made by a screwdriver, with eight replicates. Experiment 1 includes marks produced at a constant angle and direction. Experiments 2 and 3 focus on the variability between marks made at different angles and directions. Figures 4 and 5 show that the marks vary some when made at different angles and directions. We quantify these differences in the Methods and Results sections (4 and 5, respectively).
@@ 3.2 Materials
We used flat lead plates as the stratum for the screwdriver striation marks (Figure 1c). We generate the toolmarks on lead plates because we find that lead preserves fine details in the striations (Baiker et al., 2015), and it allows us to create smooth marks without high forces, which can make the slide motor stop and create jitter effects.
To scan the toolmarks in 3D, we use a handheld scanner sold by GelSight, called the Mobile 1.0X handheld instrument (Figure 1d). The GelSight instrument’s resolution (GelSight, 2023) in the $x$-$y$ (horizontal) plane is 3.45 microns, and the accuracy in the $z$ (vertical) direction is 4 microns. This scanner is hand-held and powered by a tablet, which could make it easy to use at crime scenes. It cannot be used to scan very deep marks (it has been tested up to 90 microns in depth (GelSight, 2023)) or sharp-angled surfaces that could break the gel’s surface. However, it is well-suited for scanning striation marks like the ones in this experiment and likely for many marks found at crime scenes. The scans are given by GelSight in STL format, and we converted them to x3p format, which is more common in forensic statistic analysis. The data is publicly available in [forthcoming].
@@
(b) Step 2: The black line is the location of the crosscut.
(c) Step 3: The height profile corresponding to the crosscut. The blue vertical lines are selected manually as the edges of the mark for cropping.
@@
(e) Step 5: Extracted signal. This is the residual between the black signal and the blue curve in (d). For illustration, another signal from a replicate mark is overlaid.
@@ 3.3 Signal extraction
The process to extract the digital signals from the 3D scans is shown in Figures 2a-2e. After obtaining the 3D scan from the Gelsight instrument (Figure 2a), we select a cut across the middle of the scan (Figure 2b). The reason for extracting the 2D signal from the 3D scan is that in our experimental setup, the striation marks are consistent throughout the toolmark. We checked that this was the case by taking several cross-cuts throughout the toolmark and comparing them. The information about the mark made by the screwdriver tip is contained in the relative depths of the striation marks. Going from 3D to 2D in this way is a compression that preserves the relevant information of the mark made by the screwdriver tip. The 2D signal we extract is in one way comparable to a 2D light microscope scan of the toolmark, as is used by many forensic examiners, since it contains information about the striation marks. However, our signal has the advantage of containing precise relative depth information since it is a cross-section of the toolmark.
To obtain a cross-section of the toolmark, we use the R package bulletxtrctr (Hofmann et al., 2022). We select a point in the vertical middle of the mark. We crop the edges of the profile manually (Figure 2c), and apply Gaussian smoothing to the data to model macro structures (Figure 2d), such as the plate’s shape and any scanning-specific trends. We define the signal of the screwdriver side as the difference between the profile’s height values and the fitted structure. In other words, the signal is the residual between the profile height and the fitted smoothed signal, which removes the macro structure and normalizes the signal so it is on a flat horizontal surface. Figure 2e shows the signals extracted from two different replicate marks, made by the same source.
@@ 3.4 Data
Finally, we align the signals. Figures 3-5 show examples of signals from our dataset. Figure 3 shows replicate signals from small tool 1, side A, at a fixed angle and direction. Qualitatively, the replicates look quite similar to each other. Figure 4 shows averaged replicate signals (one for each average of eight replicates), made by large tool 1 at different angles, and fixed direction. Qualitatively, the signals look similar to each other, and they look quite different from Figure 3. Figure 5 shows averaged replicate signals made by small tool 1 at different directions, and fixed angle. Note that 3 and 5 are similar because they are made with the same tool. We did not make the varying-angle marks with small tool 1 because the rig required longer tools to reach the lead plate at varying angles. It also required additional force at lower angles of attack. We leave studying these effects for future research.
@@
Figure 4: Averaged replicate signals from a single source (large tool 1), at three angles of attack (60, 70, 80), at a fixed direction (pull). Eight replicates made at each angle were averaged across angle, so only three signals are shown. The black signal is the average of the other three curves.
@@ 4.1 Method 1: Similarity matrices and clustering to studying variability by source, angle, and direction
We first cluster the signals to study the variability within tool and the variability between tools, both at a fixed setting and when varying angle and direction. This is an exploratory step that allows us to know which pairs we should consider same-source or different-source in a data-driven way. We use similarity measures that are extracted from pairwise aligned signals. We use a sliding window approach in which signals are compared in pairs by sliding one over the other, to find the lag that produces the maximum correlation between the signals.
@@ 5.1 Method 1: Similarity scores and clustering
For clustering, the first step is to calculate the similarity matrix for each experiment, meaning all the pairwise cross-correlation functions in the data. See the similarity matrices for each experiment Figures 6, 7, and 8. Qualitatively, note that there is blocking by source, angle, and direction in the similarity matrices. In all three matrices, the orange blocks off the diagonal suggest that there might be false positives if similarity is used to classify between same- and different-source. The gray blocks in the main diagonal suggest that there might be false negatives. We explore these in the next section, by means of plotting the known-match and known-non-match densities.
@@
Figure 8: Similarity matrix displayed as a heat map for the similarities between toolmarks made by 3 small screwdrivers, sides A and B, at 2 different directions (pushing and pulling), with 8 replicates each.
@@ 5.2 Method 2: Densities
Figure 9 shows the densities of similarity scores for known matches (KM) and known non-matches (KNM) for experiment 1. Note that the similarity score is the cross-correlation function (ccf), normalized to range between zero and one by adding one and dividing by two. For the KNM density, we group similarity scores by their sources and average across replicates. In other words, for the KNM density, we consider all possible combinations of 8 replicates for one source across 8 replicates for the other, resulting in 64 similarity scores for a pair of sources, and then we take the average over these 64 scores. For the KM density, we include all pairwise combinations of replicates from each source. The Beta distributions are used to estimate likelihood ratios (see Section 5.3). In Figures 10 and 11, we plot the same for experiments 2 and 3. However, we do not include the distributions or the threshold for these because they are based on a small dataset and we do not recommend classifying based on either of these experiments alone.
@@
In experiments 2 and 3 (Figures 10 and 11), we see that the densities are also separated, more clearly for experiment 2 than 3. These plots were made with much smaller datasets, so they are likely to be less smooth. Our interpretation of this separation is that the differences in angle and direction of mark generation are smaller than the differences between sources. Toolmarks made by the same source are more similar to each other than marks made by different sources, regardless of the angle and direction. This implies that the “degrees of freedom” problem is not actually so grave for toolmarks because marks made by the same source really are more similar to each other than marks made by different sources, despite the angle and direction of the tool.
@@ 5.3 Method 3: Likelihood ratio
For experiment 1, we fit parametric (Beta) distributions on each curve. The parameters for Beta distributions are chosen such that the first moment and the second moment are matched with data. Figure 9 demonstrates the two density curves along with their fitted Beta distributions. Note that the parameters for Beta distributions are derived to match the first and second moment of the data. Specifically, we obtained that for the KM curve, $\alpha=15.7494,\beta=2.0665$ and for the KNM curve, $\alpha=8.5774,\beta=11.4628$. The curves are well separated, and they intersect around a similarity score (ccf) of 0.67.
@@
For experiments 2 and 3, the middle point at which they intersect is 0.63 and 0.67, respectively. We do not fit parametric distributions to these plots or classify based on their thresholds because the densities are fit with small amounts of data, and we believed that the parametric assumptions based on these would be too strong.

### v3

@@ Abstract
Forensic toolmark analysis traditionally relies on subjective human judgment, leading to inconsistencies and lack of transparency. The multitude of variables, including angles and directions of mark generation, further complicates comparisons. To address this, we first generate a dataset of 3D toolmarks from various angles and directions using consecutively manufactured slotted screwdrivers. By using PAM clustering, we find that there is clustering by tool rather than angle or direction. Using Known Match and Known Non-Match densities, we establish thresholds for classification. Fitting Beta distributions to the densities, we allow for the derivation of likelihood ratios for new toolmark pairs. With a cross-validated sensitivity of 98% and specificity of 96%, our approach enhances the reliability of toolmark analysis. This approach is applicable to slotted screwdrivers, and for screwdrivers that are made with a similar production method. With data collection of other tools and factors, it could be applied to compare toolmarks of other types. This empirically trained, open-source solution offers forensic examiners a standardized means to objectively compare toolmarks, potentially decreasing the number of miscarriages of justice in the legal system.
@@ 1 Introduction
In this article, we present an open-source algorithm, an objective method for comparison of 3D scans of striated toolmarks. We made three contributions to the field. First, we generated three databases of toolmarks from consecutively manufactured flat-head unused screwdrivers: one to study the variability within and between tools at a fixed angle/direction, one to study the variability within and between angles of attack (80, 70, 60 with respect to the surface), and one to study the variability within and between directions of tool travel (pushing and pulling). We use the GelSight portable handheld 3D scanner, which measures the 3D topography of a solid surface using elastomeric tactile sensor technology. This dataset is available to researchers.
Second, we ran a PAM clustering algorithm Kaufman (1990) on the three databases to determine whether the similarity within source (tool-side) was higher than the similarity between sources, at a fixed angle/direction, when varying angle, and when varying direction. Then we used what we learned from the clustering step to generate the Known-Match and Known-Non-Match densities. We fit a Beta parametric distribution to these, and used an ensemble method (Guarch and Ommen, 2023), to generate likelihood ratios. Thus, a new pair or marks can be compared, and a likelihood ratio can be generated, using our method. We provide R software, an open-source implementation of this process, to make it openly and freely accessible to forensic examiners and researchers.
Third, using our method, we find that very short signals (under 1.5 mm long) cannot be compared reliably.
@@ 2 Previous work
Baiker et al. (2014) has a number of interesting contributions. They use global cross correlation to compare marks made at five angles between 15-75 degrees from the normal. Their method has high discriminatory power even at different angles of attack. They find that their automated algorithm beats human examiners in terms of false positives and the humans beat the algorithm in terms of false negatives. They also show that relying on 3D data is better than relying on 2D data, and they find evidence that although their method was trained on certain screwdrivers, it could be used to compare other tools as well. Baiker et al. (2015) find that lead preserves fine details in the striations, that it is advantageous to push the tool instead of pulling during toolmark creation for angles of attack above 45 degrees, and that toolmarks should be created as shallow as possible in the substrate material. Garcia et al. (2017) study axial rotation. Several other projects (Chumbley et al. (2009); Chumbley and Morris (2013); Grieve et al. (2014); Hadler and Morris (2018); Gambino et al. (2011); Petraco et al. (2012), to name a few) propose new machine learning and other methods to study 3D marks.
@@
From the previous work, we learn that classification is improved with 3D marks over 2D marks, and angle of attack can affect the marks. There are different classification approaches that use a variety of scores. We choose to use 3D scans, test for different angles of attack, include direction of mark generation as well, and use a simple similarity score (ccf) to ease with interpretability. There is no clustering analysis to determine what signals should be grouped together in a data-based way a priori, thus we choose to include this. Our hypothesis is that we will learn that toolmarks made at different angles and directions do cluster together, and thus there is hope for the classification of toolmarks despite being made in different ways. Furthermore, there is no likelihood ratio (LR) as far as we are aware. Our method to estimate LRs will help examiners provide their results as a numerical or verbal LR, as recommended (Willis et al., 2015).
@@ 3.1 Experimental design
Figure 1: Screwdriver tip generating a striated toolmark on the substrate material. This toolmark is made at a 50 degree angle of attack and in the “pulling” direction. Image adapted from Garcia et al. (2017).
We use a factorial design for the toolmarks generation, to allow for the study of the variability within source and between sources of marks, and how this changes as the conditions of of angle of attack and direction of mark generation change. Figure 1 shows how we define the angle (rising from the surface, so that when the screwdriver is perpendicular to the plate it is at 90 degrees) and direction of mark generation (pulling here means to the right and pushing to the left).
@@
Table 1 shows the combination of conditions for each. The factors are the angle and direction. We generate three sets of marks, which we call experiments, for a total of 560 marks. As can be seen in Table 1, we generate eight replicates per tool-side under the same conditions. Note that we used large screwdrivers for experiment 2, since varying angle required long screwdrivers to reach the lead plate. Experiments 1, 2, and 3 allow us to study the variability between marks made by the same tool, since there are eight replicates made under each condition. The angles are 60, 70, and 80 degrees with respect to the lead surface. To clarify, an angle of attack of 90 degrees means that the shaft of the screwdriver is perpendicular to the surface of the substrate. The directions are pushing and pulling, where pulling refers to a negative rake angle. For example, Figure 4 shows the marks made by a screwdriver, with eight replicates. Experiment 1 includes marks produced at a constant angle and direction. Experiments 2 and 3 focus on the variability between marks made at different angles and directions. Figures 5 and 6 show that the marks vary some when made at different angles and directions. We quantify these differences in the Methods and Results sections (4 and 5, respectively).
@@ 3.2 Materials
We used flat lead plates as the substrate for the screwdriver striation marks (Figure 2c). We generate the toolmarks on lead plates because we find that lead preserves fine details in the striations (Baiker et al., 2015), and it allows us to create smooth marks without high forces, which can make the slide motor stop and create jitter effects. We made test marks in machinable wax, copper, and aluminum. None of these materials captured the entire screwdriver mark as well as lead, so we chose lead. As this is a foundational study, it is important to observe the mark made by the entire screwdriver tip. However, future work could study how changing the material affects the marks.
To scan the toolmarks in 3D, we use a handheld scanner sold by GelSight, called the Mobile 1.0X handheld instrument (Figure 2d). The GelSight instrument’s resolution (GelSight, 2023) in the $x$-$y$ (horizontal) plane is 3.45 microns, and the accuracy in the $z$ (vertical) direction is 4 microns. This scanner is hand-held and powered by a tablet, which could make it easy to use at crime scenes. It cannot be used to scan very deep marks (it has been tested up to 90 microns in depth (GelSight, 2023) or sharp-angled surfaces that could break the gel’s surface. However, it is well-suited for scanning striation marks like the ones in this experiment and likely for many marks found at crime scenes. The scans are given by GelSight in STL format, and we converted them to x3p format, which is more common in forensic statistic analysis. The data is publicly available in [forthcoming].
@@
(b) Step 2: The black line is the location of the cross section.
(c) Step 3: The height profile corresponding to the cross section. The blue vertical lines are selected manually as the edges of the mark for cropping.
@@
(e) Step 5: Extracted signal. This is the residual between the black signal and the blue curve in (d). For illustration, another signal from a replicate mark is overlaid in the dotted profile.
@@ 3.3 Signal extraction
The process to extract the digital signals from the 3D scans is shown in Figures 3a-3e. After obtaining the 3D scan from the Gelsight instrument (Figure 3a), we select a cross section in the middle of the scan (Figure 3b). The reason for extracting the 2D signal from the 3D scan is that in our experimental setup, the striation marks are consistent throughout the toolmark. We verified this by taking several cross sections throughout the toolmark and comparing them. The information about the mark made by the screwdriver tip is contained in the relative depths of the striation marks. Going from 3D to 2D in this way is a compression that preserves the relevant information of the mark made by the screwdriver tip. The 2D signal we extract is in one way comparable to a 2D light microscope scan of the toolmark, as is used by many forensic examiners, since it contains information about the striation marks. However, our signal has the advantage of containing precise relative depth information since it is a cross-section of the toolmark.
To obtain a cross-section of the toolmark, we use the R package bulletxtrctr (Hofmann et al., 2022). We select a point in the vertical middle of the mark. We crop the edges of the profile manually (Figure 3c), and apply Gaussian smoothing to the data to model macro structures (Figure 3d), such as the plate’s shape and any scanning-specific trends. For Gaussian smoothing, we used the R loess function with span parameter set at 75%. We define the signal of the screwdriver side as the difference between the profile’s height values and the fitted structure. In other words, the signal is the residual between the profile height and the fitted smoothed signal, which removes the macro structure and normalizes the signal so it is on a flat horizontal surface. Figure 3e shows the signals extracted from two different replicate marks, made by the same source.
@@ 3.4 Data
Finally, we align the signals. Figures 4-6 show examples of signals from our dataset. Figure 4 shows replicate signals from small tool 1, side A, at a fixed angle and direction. Qualitatively, the replicates look quite similar to each other. Figure 5 shows averaged replicate signals (one for each average of eight replicates), made by large tool 1 at different angles, and fixed direction. Figure 6 shows averaged replicate signals made by small tool 1 at different directions, and fixed angle. Note that 4 and 6 are similar because they are made with the same tool. We did not make the varying-angle marks with small tool 1 because the rig required longer tools to reach the lead plate at varying angles. It also required additional force at lower angles of attack. We leave studying these effects for future research.
@@
Figure 5: Averaged replicate signals from a single source (large tool 1), at three angles of attack (60, 70, 80), at a fixed direction (pull). Eight replicates made at each angle were averaged across angle, so only three signals are shown. The black signal is the average of the other three curves. Note that these signals are wider than those of experiments 1 and 2 because they were made with larger screwdrivers.
@@ 4.1 Method 1: Similarity matrices and clustering to studying variability by source, angle, and direction
We first cluster the signals to study the variability within tool and the variability between tools, both at a fixed setting and when varying angle and direction. We use clustering as an exploratory step that allows us to know which pairs we should consider same-source or different-source in a data-driven way. In other words, we do not assume a priori that the marks made by the same tool are more similar to each other than marks made by different tools. We allow the clustering algorithm to select which sets of marks should be considered part of the same group. This step could be particularly useful in testing whether new factors, such as angle of rotation, substrate material, and force, and interactions of factors, generate marks that are so different from each other that they could be considered to have been made by a different tool.
For alignment, we use a sliding window approach (sometimes called “registration”) in which signals are compared in pairs by sliding one over the other, to find the lag that produces the maximum correlation between the signals. See Hare et al. (2017b) for a more detailed description of this process. We then use similarity measures that are extracted from pairwise aligned signals.
@@
For clustering, we use the Partitioning Around Medoids (PAM) clustering method, also called the k-medoids clustering method, which, like k-means, partitions the dataset into groups and minimizes the distance between points and their cluster centers. Unlike k-means, k-medoids selects actual data points as centers, making the clusters easier to interpret. K-medoids can also work with any dissimilarity measure, unlike k-means, which typically requires Euclidean distance. This approach is more robust to noise and outliers because it minimizes pairwise dissimilarities instead of squared Euclidean distances.
@@ 5.1 Method 1: Similarity scores and clustering
For clustering, the first step is to calculate the similarity matrix for each experiment, meaning all the pairwise cross-correlation functions in the data. See the similarity matrices for each experiment Figures 7, 8, and 9. Qualitatively, note that there is blocking by source, angle, and direction in the similarity matrices. In all three matrices, the orange blocks off the diagonal suggest that there might be false positives if similarity is used to classify between same- and different-source. The gray blocks in the main diagonal suggest that there might be false negatives. A statistical test for whether there really are “blocks” or clusters, is by running a clustering algorithm.
@@
Figure 9: Similarity matrix displayed as a heat map for the similarities between toolmarks made by 3 small screwdrivers, sides A and B, at two different directions (pushing and pulling), with 8 replicates each.
@@ 5.2 Method 2: Densities
Figure 10 shows the densities of similarity scores for known matches (KM) and known non-matches (KNM) for experiment 1. Note that the similarity score is the cross-correlation function (ccf), normalized to range between zero and one by adding one and dividing by two. For the KNM density, we group similarity scores by their sources and average across replicates. In other words, for the KNM density, we consider all possible combinations of 8 replicates for one source across 8 replicates for the other, resulting in 64 similarity scores for a pair of sources, and then we take the average over these 64 scores. For the KM density, we include all pairwise combinations of replicates from each source. The Beta distributions are used to estimate likelihood ratios (see Section 5.3). In Figures 11 and 12, we plot the same for experiments 2 and 3. However, we do not include the distributions or the threshold for these because they are based on a small dataset and we do not recommend classifying based on either of these experiments alone.
@@
In experiments 2 and 3 (Figures 11 and 12), we see that the densities are also separated, more clearly for experiment 2 than 3. These plots were made with much smaller datasets, so they are likely to be less smooth. Our interpretation of this separation is that the differences in angle and direction of mark generation are smaller than the differences between sources. Toolmarks made by the same source are more similar to each other than marks made by different sources, regardless of the angle and direction. This implies that the “degrees of freedom” problem is not actually so grave for toolmarks (within the range of 60-80 degrees) because marks made by the same source really are more similar to each other than marks made by different sources, despite the angle and direction of the tool. According to the literature (Baiker et al., 2015), the influence of angle is much larger for example in the range 80 to 110 (over the top) or 30 to 10 (very flat). Note that in Figure 8, one can observe a decrease in the correlation score when 60 is compared to 80. This trend may continue when higher angle differences are introduced. Future research could explore this trend.
@@ 5.3 Method 3: Likelihood ratio
For experiment 1, we fit parametric (Beta) distributions on each curve. Selecting Beta distributions has many advantages (Song et al., 2018). For instance, the distribution ranges from 0 to 1, which is convenient for modeling probabilities. In addition, it has two parameters, $\alpha$ and $\beta$, which allow the distribution to take on many different shapes, some of which look similar to our KM and KNM densities. The parameters for Beta distributions are chosen such that the first moment and the second moment are matched with data. Figure 10 demonstrates the two density curves along with their fitted Beta distributions. Note that the parameters for Beta distributions are derived to match the first and second moment of the data. Specifically, we obtained that for the KM curve, $\alpha=15.7494,\beta=2.0665$ and for the KNM curve, $\alpha=8.5774,\beta=11.4628$. The curves are well separated, and they intersect around a similarity score (ccf) of 0.67.
@@
We do not fit parametric distributions to the plots made with experiment 2 and 3 data, or classify based on their thresholds, because the densities are fit with small amounts of data, and we believe that the parametric assumptions based on these would be too strong. Such small datasets should not be used for statistical analysis.